Bob bungee jumped from a tower that was 350ft. Tall. It took him 3.5 seconds to reach the maximum depth of his fall, which was 50 ft. Above the ground.
If his distance from the ground varies simultaneously with time, how high above the ground was he 1.2 seconds after he jumped?

Found 2 solutions by rapture, ikleyn:
Answer by rapture(86)

This is a direct variation problem.


d = kt, where k is the constant of proportionality.


350 = k(3.5)


350/3.5 = k


100 = k


We now want to know how far into his fall he was after 1.2 seconds.


d = (1.2(100)


d = 120 feet


Bob was 120 feet after 1.2 seconds of falling.

Answer by ikleyn(52781)

.
I think that both the condition and the solution by other tutor to this problem are NONSENSE.


The formulation is fatally INACCURATE.   //   The math problems and/or Physics problems are NEVER formulated in this way.


The solution follows to wrong idea.


As formulated, the problem does not deserve any consideration.